freezing point of 4% aqueous solution of x is equal to freezing point of 12% aqueous solution of y . if molecular weight of x is a than molecular weight of y is

Grade:12th pass

1 Answers

AVNI AGRAWAL
5 years ago
we know that,
depression in freezing point= molal depression constant× molality
∂T=Kf×m                                                                                                        …....(1)
=K× wt. of solute× 1000/ Molecular wt. of solute × w t. of solvent               …....(2)
but,
%solution = 100 × wt. of solute/ wt of solvent
hence,
substituting this in equation (2)
we get,
=Kf × 10 ×%solution/ molecular wt. of solute                                                 …....(3)2
 thus,
K× 10 ×%solution(x)/ molecular wt. of solute(x) = K× 10 ×%solution(y)/ molecular wt. of solute(y)
substituting values from question we get,
4/M(x) = 12/ M(y)                      here, M denotes molecular wt.
but acc’ to question, M(x) =a
4/a=12/M(y)
rearranging it,
we get,
M(y)=3a
 
THUS, THIS IS ANSWER TO THIS QUESTION.